2t from Find the length of the parametric curve given by r(t) = 2 – logt and y(t) t=l to t= e.
Answer #1


> Ans Given X(t)= +-logt , y(t) =ąt from t= to tre We know if a= f(t), and y = g(t) then Let<ß dog B Length = de dt + .)+ ( e Length = ? Hidt ܢ ri La pool } dodol de dy debljina Jodoh hotosero) BEORD34 ons boimage) 923b v (at- +3+ (2)² dt [-(4-6)=put u= 4+1 = 0 การรับ 16+3 dt= I du 1188009 16+3 From 4+4+1=4 = 4+4=u-1 = t4 = 2+1 If t=1 » u = 40% = 1 = 5 » u = 4eft = 1 =14eH L- 4t4 dt USDdV= f = toA3 11 ) 0 0 2 ft t 4 르 S 4eH ㅗ gua 2) A4Jo ssohmul 5 retti do 2. 191 IS At AI , ㅗ 1 do - 991) s 1La Va etti २ vs 2 E ++ (lautetu +.) - toldtetti - 1) – en (15 +1) +en(6-1) ) L = 6.4947 Length = 6.4947